# Where is the Truth? The Risk of Getting Confounded in a Continual World

### v2

@@ 3.2 The ConCon Dataset
Based on the previous descriptions, we now introduce the ConCon dataset. In general, every data point is an image containing multiple objects with several attributes, building on the framework of CLEVR [18]. New data is simulated using the Blender software, and the ground truth rules and confounders are specified in JSON files. For our disjoint and strict ConCon dataset variants, we have defined 3 tasks based on a common ground truth rule, and each task has been assigned a unique confounder. The presence of the ground truth rule determines the binary class assignment of each image. An image can be either positive or negative (i.e. ground truth rule is or is not satisfied). Note that, in this continual learning setting, different tasks vary by the presence or absence of confounders, whereas the class that should be predicted remains the same (see Figure 2). Specifically, we have chosen each image to contain 4 objects, with a choice among 2 sizes, 2 materials, 3 shapes, and 8 colors. In addition, a universal unconfounded dataset contains images that are based only on the ground truth rule. Further details are provided in the appendix.
For our disjoint and strict ConCon variants the exact ground truth rule was chosen as: sphere and small cube, see Figure  3. The task-specific confounding rules are blue, metal and large, respectively.
@@
Figure 3: ConCon samples. Shown are positive and negative examples from our strict dataset. Throughout all tasks (T1, T2, T3), a sphere and a small cube (ground truth rule) must be present in a positive image but can not appear together in a negative image. Alternatively, each of the three tasks may be solved by exploiting a confounder (presence of a blue, metal, or large object, respectively). The contents of the unconfounded images depend only on the ground truth rule.
@@
In the strict setting, all positive images of a task satisfy the ground truth rule and contain the task-specific confounder. For example, for the first task, all positive images contain both a sphere and small cube and a blue object. Conversely, the negative images of a task neither satisfy the ground truth rule nor contain the task-specific confounder.
@@
In the disjoint case, the positive images of a task not only satisfy the ground truth and contain the task-specific confounder, but they further do not contain confounders specific to other tasks. Hence, for the first task all positive images contain a sphere and small cube and blue object and no metal or large object, all positive images in the second task contain a sphere and small cube and metal object and no blue or large object, all positive images in the third task contain a sphere and small cube and large object and no blue or metal object. The negative images contain neither sphere and small cube nor blue nor metal nor large, across all the tasks. All other object properties are randomly sampled.
@@
Using the above specification, we have generated 3000 train, 750 validation, and 750 test samples for both positive and negative sets of each task (i.e. 9000 images per task). The additional unconfounded dataset is generated with the same size as these splits. Overall, a total of 63000 images have thus been generated for ConCon.
@@ 4.1 Setup and Evaluation
Models. For each experiment, we use a ResNet-18 model [16] (henceforth referred to as neural network or NN) and the neuro-symbolic concept leaner (henceforth simply called NeSy) [42] that is based on the Slot Attention module [27] and Set Transformer [26]. Where the ResNet-18 model bases its decisions on learned visual features, the NeSy model first extracts object property representations before making a final prediction. We describe standard training details and hyper-parameters for both models in the Appendix.
@@
Sanity Check. To investigate whether a model is able to generalize to the unconfounded dataset after training on the confounded dataset, it must be established that the models can succeed in learning the ground truth rule when no confounders are present. To this end, we train both models only on the unconfounded dataset where the only valid solution is the ground truth rule itself. For the NN, we report an accuracy of 96.96 $\pm$ 0.44 and NeSy achieves 80.33 $\pm$ 0.91 accuracy. This confirms that the models can successfully learn the ground truth rule in principle with the right training.
@@
Comparison of Methods. We investigate nine different methods on ConCon. First, we train both models on the entire confounded dataset, where we would like to see the interplay of the confounding rules across all tasks in joint training. As a complement, we evaluate the model performance when trained on the 3 tasks one after the other in a naive sequential manner. To combat catastrophic forgetting expected in the latter we use memory-based CL methods such as ER [5], DER [4], BGS [25] as well as regularization-based methods like EWC [21] and FROMP [34] which is a mix of both. We also evaluate the performance of GDUMB [35] on ConCon that is competitive against numerous CL methods. For all the methods requiring a memory buffer, we fix the size to 100 and populate it according to what each method stipulates. Finally, we investigate a cumulative baseline, where we append the entire dataset from previous task(s) to the current one and continue training the model on the dataset accumulated so far. This is typically treated as an upper-bound to CL performance in the spirit of ER with infinite data memory [15], which will be questioned in our later findings. We provide more detailed information on the CL methods and the full training details in the Appendix.
@@
Evaluation Metrics. We measure the accuracy on the confounded test datasets across all tasks after training the model on each task. To investigate whether the model gets confounded, we further measure the accuracy on the unconfounded test set. We measure the mean accuracy calculated over 5 random seeds and report the results in Table 1. We omit standard deviations in this table, as they are negligible (we include further results on the standard deviations in the Appendix).
@@
Table 1: ConCon Evaluation. Test accuracies on disjoint and strict for NN and NeSy. Current task (T) performance is high almost throughout. The results on disjoint show that avoiding forgetting does not help with performance on unconfounded data. Strikingly, cumulative continual learning fails to find a solution on unconfounded data for strict with NN, despite joint training on all confounded data succeeding, highlighting the manifestation of insidious continual confounding.
@@
EWC [21] 100.0 99.9 100.0 50.1 60.6 47.7
FROMP [34] 100.0 99.8 99.4 99.2 97.9 50.2
ER [5] 100.0 99.6 100.0 100.0 99.6 50.1
DER [4] 100.0 99.9 100.0 99.2 98.1 50.2
BGS [25] 100.0 99.9 100.0 95.1 96.2 50.4
GDUMB [35] 64.1 62.0 69.5 63.6 61.2 48.8
@@
EWC [21] 99.6 100.0 99.9 50.2 67.2 50.2
FROMP [34] 99.6 98.9 98.4 97.9 97.3 50.5
ER [5] 99.5 99.8 99.9 99.2 99.3 50.7
DER [4] 99.6 99.7 99.9 98.6 96.9 50.9
BGS [25] 99.6 99.9 99.9 94.4 95.1 50.9
GDUMB [35] 99.4 97.8 98.5 98.0 98.2 50.3
@@
EWC [21] 100.0 99.9 100.0 49.1 48.9 50.0
FROMP [34] 100.0 59.9 69.2 67.8 59.1 51.1
ER [5] 100.0 97.7 96.4 54.6 59.6 51.0
DER [4] 100.0 98.7 96.5 52.8 55.5 49.1
BGS [25] 100.0 99.8 99.6 50.1 49.1 49.5
GDUMB [35] 92.5 62.7 53.6 57.5 54.2 50.4
@@
EWC [21] 99.9 99.9 100.0 49.1 49.4 50.1
FROMP [34] 99.8 83.4 74.7 85.1 76.3 56.2
ER [5] 99.8 99.8 98.8 68.0 83.6 61.2
DER [4] 99.7 99.4 99.5 58.8 73.2 56.5
BGS [25] 99.8 99.9 99.9 51.8 56.3 51.3
GDUMB [35] 99.4 85.6 82.7 80.1 84.2 56.7
@@
4.2 CL Methods Get Continually Confounded
We start by investigating the ConCon disjoint case, with evaluation results featured in the upper half of Table 1. Results on both NN and NeSy are reported and grouped into three categories per method: current task accuracy, old task accuracy, and accuracy on the unconfounded dataset.
For both NN and NeSy, we can see that most models successfully classify the current tasks. However, the performance of the “Naive Sequential” approach on old tasks illustrates the necessity of continual learning methods to avoid forgetting. While EWC does not appear to significantly alleviate forgetting, we can see the memory-based methods ER, DER, and BGS, as well as FROMP (benefiting mainly from the stored past examples), doing so successfully, thereby achieving close to 100% accuracy on all confounded tasks after having learned them once. GDUMB trains the model using only the buffer, resulting in bad performance for NN but is otherwise in line with the other memory-based approaches. Nevertheless, all methods on both NN and NeSy fail on the unconfounded dataset, even for the “Cumulative” setup, where all previous instances are kept when training the model sequentially. The models get continually confounded. This also remains true when all data is present in joint training, thus, explaining why avoiding forgetting is not always sufficient to avoid continual confounding.
Our results show that both models prefer learning a combination of the confounded tasks over the ground truth rule. As highlighted by our logical analysis earlier, this can be explained by our models learning the disjunction of all confounding features instead of the ground truth rule.
Ultimately, does preventing catastrophic forgetting help with continual confounding? Our evaluation on the disjoint dataset has shown that, in this setting where learning the disjunction of confounders is preferred over learning the ground truth rule, CL methods can successfully avoid forgetting. However, preventing forgetting does not help with making correct predictions on unconfounded data. In particular, the ground truth rule is also not learned when using a maximally large memory buffer (cumulative training that stores all data) or when using the NeSy approach.
@@ 4.3 CL Methods Can Get Insidiously Confounded
Consider the bottom half of Table 1. At first, it is again unsurprising that previous tasks appear to be forgotten in the naive sequential approach. The model likely learns and focuses on the new, task-specific, confounder at every stage. Possibly explained by the incompatibility of confounders (i.e. applying a rule for a task-specific confounding feature on another tasks results in no better than random prediction). EWC also shows no significant difference to the naive training setup. However, the challenge starts to unfold when taking into account memory buffer based approaches.
@@
Whereas most of these were previously able to mitigate forgetting fully across all tasks in the disjoint case, we now observe that ER, DER, and BGS perform not much better than naive sequential training in the strict scenario. Importantly, both NN and NeSy models additionally fail to make accurate predictions on the unconfounded dataset here. FROMP and GDUMB appear to sacrifice current task performance for old task performance but this does not improve performance on unconfounded data in our experiments compared to, for example, ER and gives strictly worse results than the cumulative (ER with infinite memory) approach. Contrasting this cumulative approach with joint training on all tasks reveals results that may strike the reader as surprising.
@@
In particular, we find the cumulative training with NN to seemingly focus more on the confounded tasks than on the ground truth, resulting in predictions more than 20% worse than in the joint training setup, despite training on the exact same data. The latter in fact empirically highlights how both NN and NeSy perform significantly better than random guessing when trained jointly on the strict ConCon dataset. Despite each task being confounded, the NN achieves an astonishing $95.68\%$ accuracy on the unconfounded data, compared to the achievable $96.96\%$ when trained on the unconfounded data directly (recall the initial sanity check). Whereas NeSy falls a bit shorter of its earlier baseline performance on the unconfounded dataset ($70.76\%$ vs $80.33\%$), the contrast to the observed failure on the disjoint setup is immense. In other words, we now experimentally observe our prior formalized insidious continual confounding. That is, training in a continual setup performs significantly worse than the joint setup, challenging the conventional wisdom that continual learning revolves around the accumulation of data and knowledge and questioning the role of cumulative training as a performance upper-bound [15, 32].

### v3

@@ 3.2 The ConCon Dataset
Based on the previous descriptions, we now introduce the ConCon dataset. In general, every data point is an image containing multiple objects with several attributes, building on the framework of CLEVR (Johnson et al., 2017). New data is simulated using the Blender software, and the ground truth rules and confounders are specified in JSON files. For our disjoint and strict ConCon dataset variants, we have defined 3 tasks based on a common ground truth rule, and each task has been assigned a unique confounder. The presence of the ground truth rule determines the binary class assignment of each image. An image can be either positive or negative (i.e., ground truth rule is or is not satisfied). In this continual setting, different tasks vary by the presence or absence of confounders, whereas the class that should be predicted remains the same (see Figure 2). Specifically, we have chosen each image to contain 4 objects, with a choice among 2 sizes, 2 materials, 3 shapes, and 8 colors. In addition, a universal unconfounded dataset contains images that are based only on the ground truth rule. Further details, including a dataset sheet, are provided in the appendix.
@@
Figure 3: ConCon samples. Shown are positive and negative examples from our strict dataset. Throughout all tasks (T1, T2, T3), a sphere and a small cube (ground truth rule (Refer to caption)) must be present in a positive image but can not appear together in a negative image. Alternatively, each of the three tasks may be solved by exploiting a confounder (presence of a blue, metal, or large object, respectively (Refer to caption)). The contents of the unconfounded images depend only on the ground truth rule.
@@
For our disjoint and strict ConCon variants, the exact ground truth rule was chosen as: sphere and small cube, see Figure  3. The task-specific confounding rules are blue, metal, and large, respectively. In the strict setting, all positive images of a task satisfy the ground truth rule and contain the task-specific confounder. For example, for the first task, all positive images contain both a sphere and small cube and a blue object. Conversely, the negative images of a task neither satisfy the ground truth rule nor contain the task-specific confounder.
@@
In the disjoint case, the positive images of a task not only satisfy the ground truth and contain the task-specific confounder, but they further do not contain confounders specific to other tasks. Hence, for the first task, all positive images contain a sphere and small cube and blue object and no metal or large object, all positive images in the second task contain a sphere and small cube and metal object and no blue or large object, and all positive images in the third task contain a sphere and small cube and large object and no blue or metal object. The negative images contain neither sphere and small cube nor blue nor metal nor large, across all the tasks. All other object properties are randomly sampled.
An important consideration in ConCon is the distribution shifts caused by the different confounding features and how they change across tasks. Specifically, features that act as confounders in one task can also appear as random features in other tasks. The distribution shift between the confounded tasks and the unconfounded dataset is a result of such features no longer acting as confounders, i.e., they might appear on both positive and negative images but are not informative with respect to the class label. Our synthetic data generation used for ConCon ensures minimal distribution shift between confounded and unconfounded datasets while still exhibiting confounding in one but not the other. Thus, the resulting minimal train-test mismatch is intentional and allows for investigating whether a high model accuracy is caused by the model making predictions based on the confounding features or the ground truth features.
@@
Following the specifications outlined above, we have systematically generated a comprehensive dataset comprising of 3000 train, 750 validation, and 750 test samples for both positive and negative sets of each task resulting in 9000 images per task. The additional unconfounded dataset is generated with the same size as these splits. Overall, a total of 63000 images have thus been generated for the ConCon dataset to enable thorough investigation. Precise implementation details are provided in the appendix.
In summary, ConCon is the first dataset to support the systematic investigation of the role of confounders in CL. The synthetic domains prove especially valuable in isolating and analyzing how different training mechanisms interact with data distribution characteristics. More importantly, ConCon allows us to conceive even more challenging settings, where the reader may perhaps be surprised to find that the introduction of the continual element alone is sufficient to produce confounding. And furthermore, this confounding, which was not otherwise present, even persists when sequentially accumulating the dataset with infinite memory.
@@ 4.1 Setup and Evaluation
Models. For each experiment, we use a ResNet-18 model (He et al., 2016) (henceforth referred to as neural network or NN) and the neuro-symbolic concept learner (Stammer et al., 2021) (henceforth simply called NeSy) that is based on the Slot Attention module (Locatello et al., 2020) and Set Transformer (Lee et al., 2019). Where the ResNet-18 model bases its decisions on learned visual features, the NeSy model first extracts object property representations before making a final prediction. We describe training details and hyperparameters for both models in Appendix B.
@@
Sanity Check. To investigate whether a model is able to generalize to the unconfounded dataset after training on the confounded dataset, it must be established that the models can succeed in learning the ground truth rule when no confounders are present. To this end, we train both models only on the unconfounded dataset where the only valid solution is the ground truth rule itself. For the NN, we report an accuracy of 96.96 $\pm$ 0.44 and NeSy achieves 80.33 $\pm$ 0.91 accuracy. This confirms that the models can successfully learn the ground truth rule in principle.
@@
Comparison of Methods. We investigate eleven methods on ConCon. First, we train both models on the entire confounded dataset, where we would like to see the interplay of the confounding rules across all tasks in joint training. As a complement, we evaluate the model performance when trained on the 3 tasks one after the other in a naive sequential manner. To combat catastrophic forgetting expected in the latter, we use memory-based CL methods such as ER (Chaudhry et al., 2019), DER (Buzzega et al., 2020), and BGS (Lee et al., 2024), the regularization-based method EWC (Kirkpatrick et al., 2017), and FROMP (Pan et al., 2020) which is a mix of both. We also evaluate the performance of GDUMB (Prabhu et al., 2020) on ConCon that is competitive against numerous CL methods and of PNNs (Rusu et al., 2022), where the model is expanded with each new task. We also evaluated a generative replay method (van de Ven et al., 2020), results of which can be found in Appendix D. For all the methods requiring a memory buffer, we fix the size to 100 and populate it according to what each method stipulates. Finally, we investigate a cumulative and a shuffled baseline. In the cumulative case, we append the entire dataset from previous task(s) to the current one and continue training the model on the dataset accumulated so far. This is typically treated as an upper-bound to CL performance in the spirit of ER with infinite data memory (Hayes et al., 2021), which will be questioned in our later findings. [^4] Since this is the largest possible buffer size, applying buffer-based methods like ER with larger buffer sizes is expected to yield results between what we run and the cumulative setup. To isolate the consequences of confounding, in the shuffled setup, all the data is shuffled randomly across tasks, i.e., the confounders are no longer aligned with the tasks. This breaks the temporal alignment of the confounders while maintaining the CL constraint of the cumulative setup. We provide more detailed information on the CL methods and the full training details in Appendix B.
@@
Evaluation Metrics. We measure the accuracy on the confounded test datasets across all tasks after training the model on each task. To investigate whether the model gets confounded, we further measure the accuracy on the unconfounded test set. We omit the standard deviations, as they are negligible (we include further results on the standard deviations in Appendix C).
4.2 CL Methods Get Continually Confounded
@@
Table 1: ConCon Evaluation on disjoint. Test accuracies on disjoint for NN and NeSy. Current task (T) performance is high almost throughout. Naive sequential training and EWC suffer from forgetting, showcasing a huge drop in performance on previous tasks. While other methods manage to keep high accuracies on previous tasks, our results on disjoint show that avoiding forgetting does not help with performance on unconfounded data. This holds true for both NN and NeSy.
@@
EWC (Kirkpatrick et al., 2017) 100.0 99.9 100.0 50.1 60.6 47.7
FROMP (Pan et al., 2020) 100.0 99.8 99.4 99.2 97.9 50.2
ER (Chaudhry et al., 2019) 100.0 99.6 100.0 100.0 99.6 50.1
DER (Buzzega et al., 2020) 100.0 99.9 100.0 99.2 98.1 50.2
BGS (Lee et al., 2024) 100.0 99.9 100.0 95.1 96.2 50.4
GDUMB (Prabhu et al., 2020) 64.1 62.0 69.5 63.6 61.2 48.8
PNNs (Rusu et al., 2022) 100.0 99.8 100.0 100.0 99.8 47.9
Shuffled 100.0 99.9 100.0 100.0 99.8 50.0
@@
EWC (Kirkpatrick et al., 2017) 99.6 100.0 99.9 50.2 67.2 50.2
FROMP (Pan et al., 2020) 99.6 98.9 98.4 97.9 97.3 50.5
ER (Chaudhry et al., 2019) 99.5 99.8 99.9 99.2 99.3 50.7
DER (Buzzega et al., 2020) 99.6 99.7 99.9 98.6 96.9 50.9
BGS (Lee et al., 2024) 99.6 99.9 99.9 94.4 95.1 50.9
GDUMB (Prabhu et al., 2020) 99.4 97.8 98.5 98.0 98.2 50.3
PNNs (Rusu et al., 2022) 99.5 100.0 100.0 99.6 100.0 49.6
Shuffled 99.8 99.7 99.9 99.7 99.9 50.8
@@
We start by investigating the ConCon disjoint case, with evaluation results featured in the Table 1. Results on both NN and NeSy are reported and grouped into three categories per method: current task accuracy, old task accuracy, and accuracy on the unconfounded dataset.
For both NN and NeSy, we can see that most models successfully classify the current tasks. However, the performance of the “Naive Sequential” approach on old tasks illustrates the necessity of continual learning methods to avoid forgetting. While EWC does not appear to significantly alleviate forgetting, we can see the memory-based methods ER, DER, BGS, and FROMP (benefiting mainly from the stored past examples), as well as PNNs (benefiting from additional classification heads) doing so successfully, thereby achieving close to 100% accuracy on all confounded tasks after having learned them once. GDUMB trains the model using only the buffer, resulting in bad performance for NN but is otherwise in line with the other memory-based approaches. Nevertheless, all methods on both NN and NeSy fail on the unconfounded dataset, even for the “Cumulative” setup, where all previous instances are kept when training the model sequentially. The models get continually confounded. This also remains true in both cases where a model is trained on data from all tasks at the same time, no matter if it is trained continually (shuffled) or on the full dataset at the same time (joint). Since even the joint training setup does not learn the ground truth rule, it is unsurprising that the continual learning methods fail to do so as well.
Our results show that both models prefer learning a combination of the confounded tasks over the ground truth rule, as they provide reliable shortcuts for solving individual tasks during sequential training. As highlighted by our logical analysis, this can be explained by our models learning the disjunction of all confounding features instead of the ground truth rule, despite the models’ general ability to learn the ground truth rule, as previously shown in our sanity check.
Table 2: ConCon Evaluation on strict. Test accuracies on strict for NN and NeSy. All continual learning methods fail to achieve high accuracies on previous tasks, especially for NN. Strikingly, even cumulative continual learning fails to find a solution on unconfounded data for strict with NN, despite joint training on all confounded data succeeding. While shuffled training also sees a small drop in accuracy, a big gap to joint training remains. This shows that the majority of the difference between joint and cumulative training is not caused by the CL setting, but by the confounding setup, highlighting the manifestation of insidious continual confounding.
Current Task Acc. Old Task Acc. Unconf.
Method T1 T2 T3 T1@T3 T2@T3
@@
EWC (Kirkpatrick et al., 2017) 100.0 99.9 100.0 49.1 48.9 50.0
FROMP (Pan et al., 2020) 100.0 59.9 69.2 67.8 59.1 51.1
ER (Chaudhry et al., 2019) 100.0 97.7 96.4 54.6 59.6 51.0
DER (Buzzega et al., 2020) 100.0 98.7 96.5 52.8 55.5 49.1
BGS (Lee et al., 2024) 100.0 99.8 99.6 50.1 49.1 49.5
GDUMB (Prabhu et al., 2020) 92.5 62.7 53.6 57.5 54.2 50.4
PNNs (Rusu et al., 2022) 100.0 99.8 99.8 100.0 99.8 49.2
Shuffled 79.9 93.4 95.3 95.2 96.7 89.5
@@
EWC (Kirkpatrick et al., 2017) 99.9 99.9 100.0 49.1 49.4 50.1
FROMP (Pan et al., 2020) 99.8 83.4 74.7 85.1 76.3 56.2
ER (Chaudhry et al., 2019) 99.8 99.8 98.8 68.0 83.6 61.2
DER (Buzzega et al., 2020) 99.7 99.4 99.5 58.8 73.2 56.5
BGS (Lee et al., 2024) 99.8 99.9 99.9 51.8 56.3 51.3
GDUMB (Prabhu et al., 2020) 99.4 85.6 82.7 80.1 84.2 56.7
PNNs (Rusu et al., 2022) 99.8 100.0 100.0 99.8 100.0 49.9
Shuffled 87.8 92.5 89.2 88.2 92.6 70.1
@@
Ultimately, does preventing catastrophic forgetting help with continual confounding? Our evaluation on the disjoint dataset showed that, in this setting where learning the disjunction of confounders is preferred over learning the ground truth rule, CL methods can successfully avoid forgetting. However, preventing forgetting does not help with making correct predictions on unconfounded data. In particular, the ground truth rule is also not learned when using a maximally large memory buffer (cumulative training that stores all data) or when using the NeSy approach.
@@ 4.3 CL Methods Can Get Insidiously Confounded
The experimental results on the strict dataset are shown in Table 2. First, consider the only approach that outputs different predictions depending on which data the task is coming from: PNNs. Here, the model has different classification heads for each task, enabling it to fully avoid forgetting and perform well on current and past tasks. However, this leads to learning the respective tasks individually, failing to identify the underlying ground truth rule that applies to all tasks. Considering the more standard continual learning setups, it is again unsurprising that previous tasks appear to be forgotten in the naive sequential approach. The model likely learns and focuses on the new, task-specific confounder at every stage. This may be explained by the incompatibility of confounders (i.e., applying a rule for a task-specific confounding feature on another task results in no better than random prediction). EWC also shows no significant difference to the naive training setup. However, differences between our dataset variants start to show when considering the memory-based approaches.
@@
Whereas most of these were previously able to mitigate forgetting fully across all tasks in the disjoint case, we now observe that ER, DER, and BGS perform not much better than naive sequential training in the strict scenario. Importantly, both NN and NeSy models additionally fail to make accurate predictions on the unconfounded dataset. FROMP and GDUMB appear to sacrifice current task performance for old task performance, but this does not improve performance on unconfounded data in our experiments compared to, for example, ER and gives strictly worse results than the cumulative (ER with infinite memory) approach. All standard CL methods thus not only fail to perform well on the unconfounded data but also do not get close to successfully avoiding forgetting, with accuracies on old tasks not even exceeding 60% except for FROMP, where current task performance is much lower. This shows how confounded datasets can be especially challenging for CL methods, even when those successfully avoid forgetting in easier problems. While the cumulative approach, benefiting from its infinite buffer size, performs much better in comparison, contrasting it with joint and shuffled training on all tasks reveals results that may strike the reader as surprising.
@@
For NeSy, cumulative training and joint training perform similarly well, but they both do not reach their earlier baseline performance on the unconfounded dataset ($70.76\%$ vs $80.33\%$). More interestingly, we find the cumulative training with NN to seemingly focus more on the confounded tasks than on the ground truth, resulting in predictions more than $20\%$ worse than in the joint training setup, despite training on the exact same data. The latter in fact empirically highlights how both NN and NeSy perform significantly better than random guessing when trained jointly on the strict dataset. Despite each task being confounded, the NN achieves an astonishing $95.68\%$ accuracy on the unconfounded data, compared to the achievable $96.96\%$ when trained on the unconfounded data directly (recall the initial sanity check). While some of the differences between joint and cumulative training may not be solely explained by the presence of confounders, but also by the continual learning setup itself, we observe in our shuffled experiment that training continually on the same data yields substantially better results when the tasks are not aligned with the confounding features. In other words, we now experimentally observe our prior formalized insidious continual confounding. That is, training in a continual setup performs significantly worse than the joint setup, challenging the conventional wisdom that continual learning revolves around the accumulation of data and knowledge and questioning the role of cumulative training as a performance upper-bound (Hayes et al., 2021; Mundt et al., 2023b).